OpenStudy (anonymous):

Im having some math trouble. I know this is pretty basic, but I completely forgot what the formula or whatever is called so I cant look it up in my book. If anyone could help me with this equation I would really appreciate it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For an example, lets say the number of drops is 60

OpenStudy (anonymous):

go on google and they have a conversion cal that will do it for you.

OpenStudy (anonymous):

No, I need to do it this way. This is what my teacher wants. I need to show my work.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So youre not going to help?

OpenStudy (anonymous):

i will try to help if i knew how to do it.

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

Ratios with an unknown. Let me start with a simpler case. One with just 2 types of units. Then expand this to three. For example: 1 US gallon = 128 US ounces. So I can divide either side by the other and get 1. \(\dfrac{1 gal}{128 oz}=1\) \(\dfrac{128 oz}{1 gal}=1\) So, lets say I want to convert a number of oz to gallons. Then I need to set up the ratio in such a way that the oz cancel. Let me do this with 200 ounces. \(\dfrac{1 gal}{128 oz}= \dfrac{?}{200 oz}\implies \) \(\dfrac{1 gal}{128 oz}\cdot \dfrac{200 oz}{1}= \dfrac{?}{200 oz}\cdot \dfrac{200 oz}{1}\implies \) \(\dfrac{1 gal}{128\color{red}{\cancel{ oz}}}\cdot \dfrac{200 \color{red}{\cancel{ oz}}}{1}= \dfrac{?}{\color{red}{\cancel{200 oz}}}\cdot \dfrac{\color{red}{\cancel{200 oz}}}{1}\implies \) \(\dfrac{ 200 gal}{128 }= \dfrac{?}{1} \implies\) \(\dfrac{ 25 gal}{16 }= ? \implies\) \(1.5625 gal = ?\) So 200 oz is 1.5625 gal.

OpenStudy (anonymous):

Okay so 60 drops with my equation would be 2.956 mL?

OpenStudy (e.mccormick):

Well, with a drops to mL converter, that is what I get. I would need to know the number of drops in a tsp to show it any other way. Basically, you use anything that = 1 as a basis to do the conversion.

OpenStudy (anonymous):

64 drops per tsp

OpenStudy (anonymous):

eWhich is about 13 drops per mL

OpenStudy (e.mccormick):

OK. So 64 drops per tsp and 4.93 ml per tsp. Then you use that information to say it is also 64 drops per 4.93 ml. Then it solves like the galons to oz.

OpenStudy (anonymous):

So I was right with what I got?

OpenStudy (e.mccormick):

Lets see.

OpenStudy (e.mccormick):

This would be the setup at the start: \(\dfrac{4.93mL}{1 tsp}\cdot \dfrac{1 tsp}{64 drops} = \dfrac{?}{60 drops}\)

OpenStudy (e.mccormick):

\(\dfrac{4.93mL}{\cancel{1 tsp}}\cdot \dfrac{\cancel{1 tsp}}{64 drops}\cdot\dfrac{60 drops}{1} = \dfrac{?}{60 drops}\cdot\dfrac{60 drops}{1}\) \(\dfrac{4.93mL}{64 \cancel{drops}} \cdot\dfrac{60 \cancel{drops}}{1} = \dfrac{?}{\cancel{60 drops}}\cdot\dfrac{\cancel{60 drops}}{1}\)

OpenStudy (anonymous):

I get 5.25 and that doesnt seem right

OpenStudy (e.mccormick):

\(\dfrac{60 \cdot 4.93mL}{64 } = ?\)

OpenStudy (anonymous):

4.62...but again that doesnt seem right because thats not how the equation is set up

OpenStudy (e.mccormick):

Well, for 60 drops, it should be very close to 1 tsp if there are 64 in a tsp. 60/64 = .9375, or 93.75%! That is almost a full tsp.

OpenStudy (e.mccormick):

So make sure that 64 drops in a teaspoon is right.

OpenStudy (e.mccormick):

If I recall, the drops per volume is dependent on the viscosity of the fluid. So if you are given that a particular fluid is 64 drops per oz, that is what you use. For some thick fluid, it could be that. For water is is like 120.